OpenStudy (anonymous):

A certain radioactive element has a half-life of 1 h. If you start with a 1-g sample of the element at noon, how much of this same element will be left at 3:00 PM? (a) 0.5 g (b) 0.25 g (c) 0.125 g (d) 0.0625 g @Abhisar

OpenStudy (abhisar):

Hello @beccamarx19 !

OpenStudy (abhisar):

Hmm noon means 12 PM, so 3 half lifes have elapsed.

OpenStudy (abhisar):

I taught u the formula for calculating the % remaining earlier. Using that one we can calculate the answer

OpenStudy (abhisar):

Here, n= 3 \(\huge\frac{1}{2^3}\) = 1/8 of the original amount i.e 1/8 of 1g = 0.125g

OpenStudy (unklerhaukus):



OpenStudy (unklerhaukus):